# An experimental platform for levitated mechanics in space

### v1

@@ 4 Payload Design
One of the payload’s end-plates was designed as a heat sync for the LEGO electronics board and FPGA (see Section 5), while the other was recessed to hold the optical trap’s photodetector (see Section 6). Both end-plates were bolted to the stainless steel baseplate which had was designed with M10 through-holes for bolting the payload to the Nyx capsule. The end-plates were also held together at the top of the payload by aluminium stringers, while larger stringers between the two end-plates held the vacuum chamber near the payload’s midpoint. Another stringer was added between the top stringers to hold the optical fibre mating sleeves. Stainless steel cover-plates were attached to the stringers and end-plates to enclose the payload. Images of the completed payload can be seen in Fig. 3. To ensure that all structural components would be secure throughout the mission, all threaded parts used locking threads (Spiralock) and a threadlocking agent (Loctite 242). Similarly, all internal wires and optical fibres were secured with dots of epoxy (3M DP190) at least every 2 cm.
@@ 5 Electronics
The laser (Thorlabs DFB1550P) used in the experiments is driven by a custom ‘LEGO’ electronics board, produced by a joint team from Twin Paradox Labs and Surrey Space Centre [40, 41], that was controlled by an FPGA (Opal Kelly XEM7310-A200). The LEGO board enables measurement and control of the laser’s output and thermo-electric controller, and also has capabilities for analogue input and output, PID control and 16-bit data saving. The LEGO board and FPGA jointly draws approximately 7 W.
The 28 VDC supply from the Nyx capsule is passed through a fifth-order low pass filter to protect both the Nyx capsulse and the payload from voltage fluctuations. The input is then internally converted to a 5 VDC line and $\pm$12 VDC rails. The ±12 VDC is used to power the optical trap’s photodetector (Thorlabs PDBEVAL1) while the 5 VDC powers the remaining systems, namely the LEGO board and attached FPGA, the optical trap’s piezo kicking system and a microcontroller (Arduino Portenta H7) used to control internal signals and acquire experimental data. The 5 VDC also provides a reverse bias to the magnetic trap’s photodetector (Thorlabs FDG05) to increase the detector’s sensitivity.

### v2

@@ 4 Payload Design
One of the payload’s end-plates was designed as a heat sync for the laser driver PCB and FPGA (see Section 5), while the other was recessed to hold the optical trap’s photodetector (see Section 6). Both end-plates were bolted to the stainless steel baseplate which had was designed with M10 through-holes for bolting the payload to the Nyx capsule. The end-plates were also held together at the top of the payload by aluminium stringers, while larger stringers between the two end-plates held the vacuum chamber near the payload’s midpoint. Another stringer was added between the top stringers to hold the optical fibre mating sleeves. Stainless steel cover-plates were attached to the stringers and end-plates to enclose the payload. Images of the completed payload can be seen in Fig. 3. To ensure that all structural components would be secure throughout the mission, all threaded parts used locking threads (Spiralock) and a threadlocking agent (Loctite 242). Similarly, all internal wires and optical fibres were secured with dots of epoxy (3M DP190) at least every 2 cm.
@@ 5 Electronics
The laser (Thorlabs DFB1550P) used in the experiments is driven by a custom laser driver PCB, produced by a joint team from Twin Paradox Labs and Surrey Space Centre [40, 41], that was controlled by an FPGA (Xilinx Artix 7 A200). The laser driver enables measurement and control of the laser’s output and thermo-electric controller, and also has capabilities for analogue input and output, PID control and 16-bit data saving. The laser driver and FPGA jointly draws approximately 7 W.
The 28 VDC supply from the Nyx capsule is passed through a fifth-order low pass filter to protect both the Nyx capsulse and the payload from voltage fluctuations. The input is then internally converted to a 5 VDC line and $\pm$12 VDC rails. The ±12 VDC is used to power the optical trap’s photodetector (Thorlabs PDBEVAL1) while the 5 VDC powers the remaining systems, namely the laser driver PCB and attached FPGA, the optical trap’s piezo kicking system and a microcontroller (Arduino Portenta H7) used to control internal signals and acquire experimental data. The 5 VDC also provides a reverse bias to the magnetic trap’s photodetector (Thorlabs FDG05) to increase the detector’s sensitivity.